OpenStudy (anonymous):

need help please... The length of a rectangular jewelry box is 3 inches more than twice the width. the perimeter is 18 inches. find the length and width so far I know how to do this part: 18=2(2x+3)+2x 18=4x+6+2x

OpenStudy (anonymous):

L=12?

OpenStudy (anonymous):

7 and 2 x = 2 2x + 3 = 7

OpenStudy (anonymous):

awe gash i typed it wrong on my hw i got it wrong grr=/

OpenStudy (anonymous):

your equations are correct btw

OpenStudy (anonymous):

The length of a rectangular jewelry box is 5 inches more than twice the width. the perimeter is 28 inches. find the length and width

OpenStudy (anonymous):

thanks:) i love math but if i understand it its better

OpenStudy (anonymous):

28=2(2x+5)+2x 28=4x+10+2x

OpenStudy (anonymous):

am I right so far?

OpenStudy (anonymous):

28=6x+10 -10 -10 18=6x x=3

OpenStudy (anonymous):

yep its right

OpenStudy (anonymous):

so is that length or width cuz i have to find both=/

OpenStudy (anonymous):

u made ur equation urself so u should know better :P in ur case it's width

OpenStudy (anonymous):

hahaha im sucha dumbblonde

OpenStudy (anonymous):

ok if that width i have to find the length now

OpenStudy (anonymous):

yeah length is 5 cm longer than double width :D

OpenStudy (anonymous):

double the width so 9?

OpenStudy (anonymous):

i wanna finish this so i can slp away..

OpenStudy (anonymous):

am I right or wrong for the answer

OpenStudy (anonymous):

err i mean not double, twice

OpenStudy (anonymous):

soo 10?

OpenStudy (anonymous):

twice the width and 5 more

OpenStudy (anonymous):

how u got 10?

OpenStudy (anonymous):

lol u said twice the width & more

OpenStudy (anonymous):

5inches more and you shouldn't listen to me, you should read ur problem :D

OpenStudy (anonymous):

oh goodness....

OpenStudy (anonymous):

oh ur so funny you! 0_0

OpenStudy (anonymous):

please help me finish=/

OpenStudy (anonymous):

the energy drink i had is knockin me out lol

OpenStudy (anonymous):

double width and add 5 :D

OpenStudy (anonymous):

so width is 5?

OpenStudy (anonymous):

no u wrote that x=3

OpenStudy (anonymous):

lmao oh goodness see

OpenStudy (anonymous):

so 3(3)=15

OpenStudy (anonymous):

lol im confusing myself

OpenStudy (anonymous):

double means multiply by 2

OpenStudy (anonymous):

3 x 2? 6

OpenStudy (anonymous):

and add 5

OpenStudy (anonymous):

11

OpenStudy (anonymous):

WOOOOO thats correct answer@@@

OpenStudy (anonymous):

yayyyyyyyyyyyyyy:D

OpenStudy (anonymous):

go tomas! & go me lmao xD

OpenStudy (anonymous):

ding ding ding i finally got it

OpenStudy (nubeer):

i think Length = 7 and width= 2

OpenStudy (anonymous):

nope that was the problem be4... thaanks anyway

OpenStudy (nubeer):

oh ok

OpenStudy (anonymous):

no u are incorrect 2*2+7*2=18 and it should be 28

OpenStudy (anonymous):

oh wow i killed it... o_0

OpenStudy (anonymous):

sry tomas

OpenStudy (anonymous):

nubeer is incorrect

OpenStudy (anonymous):

you are correct :D

OpenStudy (anonymous):

I'm correct ;) 100%

OpenStudy (nubeer):

sorry i ddnt read it was equal to 28. .. well never mind. :P

OpenStudy (anonymous):

nice try tho